A psychologist wanted to assess the relationship between depression and self-esteem. She believes that as depression increases, self-esteem decreases. She asked a group of 25 participants to complete both Beck’s Depression Inventory (Beck, Ward, Mendelson, Mock, & Erbaugh, 1961) and Rosenberg’s Self-Esteem Scale (1965). The data for each participant is as follows:

Depression (X)

Self-Esteem (Y)

60

11

27

34

15

33

30

24

56

14

52

16

40

23

49

33

24

27

38

26

48

17

58

10

22

39

51

18

41

22

56

17

49

16

54

13

52

12

41

23

35

28

20

35

55

15

16

32

28

20

Descriptive Statistics

For Depression:

  1. What is the scale of measurement of the variable?
  2. Is the distribution normally distributed, positively skewed, or negatively skewed?
  3. Based on your answers two the first two questions, what would be the best statistics to use to describe this variable (central tendency and variability)?
  4. Why would these be the correct/best choice?
  5. What are the values of these statistics (the appropriate statistics ONLY – not all)?

For Self-Esteem:

  1. What is the scale of measurement of the variable?
  2. Is the distribution normally distributed, positively skewed, or negatively skewed?
  3. Based on your answers two the first two questions, what would be the best statistics to use to describe this variable (central tendency and variability)?
  4. Why would these be the correct/best choice?
  5. What are the values of these statistics (the appropriate statistics ONLY – not all)?

Correlation

  1. What is the predictor variable?
  2. What is the criterion variable?
  3. Correlation: Use the steps of hypothesis testing to evaluate this hypothesis.
    1. State the hypotheses.
    2. Determine the critical value of your statistic for rejecting the null based on the given alpha level/sample size.
    3. What is the value of the statistic.
    4. Based on this statistic alone, would you reject or accept the null hypothesis?
  4. What is the effect size? What does this mean?

Regression

  1. What is the slope of the regression line?
  2. What is the intercept of the regression line.
  3. Insert the calculated values for the slope and intercept into the regression line equation and enter that here.
  4. Determine the predicted self-esteem scores for participants who have the following levels of depression:
    1. 28: Y’ =
    2. 42: Y’ =
    3. 54: Y’ =
Answer #1

For dipression

i) The scale of measurement of the variable: Quantitative data (numerical values are provided).

ii) The data is normally distributed. We plot normal probability in Minitab.

iii) Since we have numerical measurements we can calculate mean and variance (standard deviation).

iv) Since the data is numerical and normally distributed, the mean and variance are the correct measures to calculate.

v) Mean=sum of all observations/total no of observation=1017/25=40.68

Variance=(sum of all squared observations/total no of observation)-(mean)^2=196.2176

For Self-esteem

i) The scale of measurement of the variable: Quantitative data (numerical values are provided).

ii) The data is normally distributed. We plot normal probability in Minitab.

iii) Since we have numerical measurements we can calculate mean and variance (standard deviation).

iv) Since the data is numerical and normally distributed, the mean and variance are the correct measures to calculate.

v) Mean=sum of all observations/total no of observation=558/25=22.32

Variance=(sum of all squared observations/total no of observation)-(mean)^2=69.0176

Correlation

i) Predictor variable: Self-esteem

ii) Criterion variable: Dipression

X^2 Y^2 XY
3600 121 660
729 1156 918
225 1089 495
900 576 720
3136 196 784
2704 256 832
1600 529 920
2401 1089 1617
576 729 648
1444 676 988
2304 289 816
3364 100 580
484 1521 858
2601 324 918
1681 484 902
3136 289 952
2401 256 784
2916 169 702
2704 144 624
1681 529 943
1225 784 980
400 1225 700
3025 225 825
256 1024 512
784 400 560
Total 46277 14180 20238
X Y
Mean 40.68 22.32
Varince 196.2176 69.0176
Covariance -98.4576
Correlation -0.84606

Null hypothesis H0: ρ = 0
Alternative hypothesis HA: ρ ≠ 0 or HA: ρ < 0 or HA: ρ > 0

Test statistic: t=r\sqrt{n-1}/\sqrt{(1-r^2))} , where r is correlation and n is total no of observation

t=-14.5849

|t|=14.5849

p-value=0.0000

Since the p-value< 5%.

We may reject H0.

Regression

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.846058
R Square 0.715815
Adjusted R Square 0.703459
Standard Error 4.617286
Observations 25
ANOVA
df SS MS F Significance F
Regression 1 1235.096 1235.096 57.93314 9.97E-08
Residual 23 490.3445 21.31933
Total 24 1725.44
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 42.73231 2.836353 15.06593 2.09E-13 36.86487 48.59976 36.86487 48.59976
X Variable 1 -0.50178 0.065925 -7.61138 9.97E-08 -0.63815 -0.3654 -0.63815 -0.3654

Here intercept: b0=42.73231

slope: b1= -0.50178

a) x=28: Y=28.68254

b) x=42: Y=21.6576

c) x=54: Y=15.6363